# Rethinking Graph Structure Learning in the Era of LLMs

### v4

@@ F.1 Topological Information Capturing Capability of the Encoding Tree
Lemma 1 (Locality and Additivity of Structural Entropy[12]).
@@ F.2 Global information constraints of the low-level communities
the entropy term of the low-level community $\mathcal{C}^{\ell}_{\phi}$ becomes more localized [12].
@@ F.3 The Error Bound Between Soft Labels and True Labels
The proof leverages three key lemmas derived from the Transformer architecture in [25].
@@
Let $t^{\star}_{\alpha}$ be the aggregated text of node $\alpha$ using threshold $\epsilon$ (Eq. 2). The raw soft label error $\Delta_{\alpha}^{\text{raw}}=|y^{cls}_{\alpha}-y_{\alpha}$| satisfies:
@@
the threshold $\epsilon$ filters neighbors with $w_{\alpha\beta}\geq\epsilon$ (Eq. 2 and Eq. 3). The multi-head attention reinforces local semantic consistency across subspaces, bounding the error proportionally to $(1-\epsilon)$ [25]. ∎
@@
The residual structure $\text{LayerNorm}(x+\text{Sublayer}(x))$ suppresses feature drift, with $d_{\text{model}}$ as the hidden dimension [25]:
@@
The learning rate design $lrate=d_{\text{model}}^{-0.5}\cdot\min(\cdot)$ indicates that increasing the model dimensionality necessitates a reduction in learning rate to preserve training stability. A larger $d_{\text{model}}$ enhances the model’s feature representation capacity, while the error decay rate per unit dimension scales proportionally with $d_{\text{model}}^{-0.5}$ [25]. ∎
@@ Appendix G Datasets
We evaluate LLaTA and baselines on 11 widely adopted TAG datasets across multiple domains, including 4 citation networks [3], 1 knowledge network [3], 2 social networks [9], and 4 e-commerce networks [30]. Among these, the Ratings and Child exhibit heterophily, while the remaining datasets follow homophily. The details of these TAG datasets are shown in Table 6. The description of the datasets for each domain is as follows:
@@
  • Citation Networks  [3]. Cora, Citeseer, Pubmed and ArXiv are benchmark datasets of citation networks. Nodes represent paper, and edges represent citation relationships. The features are obtained through pre-trained language model, and labels denote their academic fields.
  • Knowledge Networks  [3, 19]. WikiCS is a benchmark dataset of knowledge networks. Nodes represent articles in the field of computer science, and edges represent hyperlinks between these articles. The features are obtained through pre-trained language model, and labels denote different branches of computer science.
  • Social Networks  [9]. Instagram and Reddit are benchmark datasets of social netowrks. Nodes represent users, and edges represent social relationhships. The features are obtained through pre-trained language model, and labels correspond to different types of users.
  • E-commerce Networks  [30]. Ratings, Child, History and Photo are benchmark datasets of e-commerce networks. Nodes represent products, and edges represent relationships such as co-purchase or co-view. The features are obtained through pre-trained language model, and labels correspond to product categories or user ratings.
@@ Appendix H Baselines
In this section, we provide a brief description for each baseline used in the experiments. It includes 5 coupled GSL methods [2, 6, 36, 28, 42], 6 decoupled GSL methods [10, 17, 13, 15, 20, 1], and 3 LLM-based GSL methods[7, 34]. Based on this, we apply prevalent GNN (GCN [11], GAT [26], GraphSAGE [8]) and mainstream LLM-GNN (GLEM [35], ENGINE [41]) as the downstream backbone.
@@
  • GCN [11]. GCN is among the most widely adopted GNN architectures, as it introduces a first-order approximation of a localized spectral filter for graph-structured data.
@@
  • GAT [26]. GAT incorporates a self-attention mechanism to compute importance scores for different neighboring nodes, enabling more effective aggregation of neighborhood information.
@@
  • GraphSAGE [8]. GraphSAGE [14] is an inductive framework that generates node embeddings by leveraging node features and using a sampling-based approach to aggregate information from the local neighborhood, enabling scalable representation learning.
@@
  • GLEM [35]. GLEM proposes a scalable framework that integrates LLMs with GNNs through a variational EM algorithm. By alternating between optimizing LMs for textual semantics and GNNs for structural information, GLEM enables mutual distillation without costly end-to-end training, achieving promising performance on large-scale text-attributed graphs.
@@
  • ENGINE [41]. ENGINE introduces an LLM-enhanced graph learning framework that combines LLMs with GNNs through a mutual integration mechanism. It leverages LLMs to enrich node representations with semantic knowledge while using GNNs to capture structural dependencies, and aligns the two via iterative refinement. This design improves both effectiveness and robustness on text-attributed graphs, showing promising performance across multiple benchmarks.
@@
  • IDGL [2]. IDGL is a framework that learns a refined graph structure by leveraging cosine similarity of features and top-k threshold refinement, combining it with the original graph for joint optimization.
@@
  • SLAPS [6]. SLAPS is a graph learning framework that uses MLP-based embeddings and kNN to build the graph. The adjacency matrix is symmetrized and normalized, and a self-supervised denoising autoencoder is used to update both the graph and model parameters.
@@
  • GAUGO [36]. GAUGO uses a graph auto-encoder to learn a new graph structure. It predicts edges based on node features, refines the edge probabilities with Gumbel sampling, and fuses the generated graph with the original one before training. Both the generated graph and model parameters are updated during optimization.
@@
  • HESGSL [28]. HESGSL uses hierarchical embeddings to capture both local and global graph structures. It employs self-supervised learning with clustering and attention mechanisms to enhance feature representation and performance on downstream tasks.
@@
  • SEGSL [42]. SEGSL constructs a graph by fusing a kNN graph with the original graph, leveraging structural entropy and encoding tree to refine edges and hierarchically extract community structures. The graph and model parameters are optimized jointly during training.
@@
  • ProGNN [10]. ProGNN refines graph structures by enforcing low-rank, sparsity, and similarity constraints, jointly optimizing the learned graph and model parameters without predefined GSL bases or graph fusion.
@@
  • SUBLIME [17]. SUBLIME constructs GSL graphs using an anchor view (original graph) and a learner view. The learn view is initialized with kNN and optimized with parameterized or non-parameterized methods, followed by post-processing steps like top-k filtering and symmetrization. Contrastive learning between the two views refines the graph for downstream tasks.
@@
  • STABLE [13]. STABLE employs Graph Contrastive Learning (GCL) to generate robust graphs by perturbing node similarities and edges. Positive samples use slight perturbations, while negative samples use shuffled features. The graph refinement step applies a top-k filtering strategy on the node similarity matrix to retain helpful edges while removing adversarial ones.
@@
  • CoGSL [15]. CoGSL constructs GSL graphs by combining the original graph with additional views, such as adjacency or kNN graphs. Node embeddings are generated via GCNs, and connection probabilities are estimated and refined using InfoNCE loss to maximize mutual information. The final graph is obtained through early fusion and updated with model parameters.
@@
  • BORF [20]. BORF is a curvature-based graph rewiring approach aimed at addressing over-smoothing and over-squashing in GNNs. It leverages Ollivier-Ricci curvature to identify edges with extreme curvature values—positive for over-smoothing and negative for over-squashing. This process preserves the graph’s topology while improving message-passing efficiency and enhancing GNN performance on downstream tasks.
@@
  • DHGR [1]. DHGR enhances GNN performance on heterophily graphs by preprocessing the graph structure through rewiring. It systematically adds homophilic edges and removes heterophilic ones based on label and feature similarities, improving node classification accuracy, particularly for heterophily settings.
@@
  • GraphEdit [7]. GraphEdit refines graph structures by integrating LLMs with a lightweight edge predictor to enhance graph representation learning. By leveraging instruction-tuning and training edge predictor, GraphEdit effectively denoises noisy connections and identifies implicit relationships among nodes.
@@
  • LLM4RGNN [34]. LLM4RGNN is a framework that enhances the adversarial robustness of GNNs using LLMs. It detects malicious edges and restores critical ones through a purification strategy, leveraging fine-tuned LLMs for edge prediction and correction. This approach effectively mitigates topology attacks, improving GNN performance across attack scenarios.
@@
  • LangGSL [24]. LangGSL presents a novel framework that integrates LLMs with GSL in an end-to-end manner. It simultaneously optimizes node features and graph connectivity by using LLMs to denoise textual attributes while GSL modules refine structural relations. This joint design enhances both semantic understanding and topological quality, leading to promising performance across diverse text-attributed graph datasets.
@@ Appendix I Experimental Setup
Current graph structure learning (GSL) methods face significant efficiency bottlenecks, particularly in large-scale or noisy graphs. Traditional approaches—including metric learning, probabilistic modeling, and direct optimization—often require computationally expensive pairwise similarity calculations (e.g., kernel functions or attention mechanisms) or iterative optimization of adjacency matrices[40]. These operations exhibit quadratic or higher complexity relative to node counts, limiting scalability. For LLM-based GSL Methods (e.g., GraphEdit [7], LLM4RGNN [34]), efficiency issues are further exacerbated by the inherent overhead of large-model inference and fine-tuning. Empirical results reveal that these methods exceed 72 hours on larger datasets among the 11 benchmarks (As shown in Table 1), primarily due to their reliance on full-parameter tuning and dense graph operations.
Compared to existing paradigms, our framework achieves superior efficiency by: (a) Topology-Aware Context Enhancement: The Community-of-Thought mechanism provides high-quality topological context to guide LLM reasoning, eliminating the need for fine-tuning. By leveraging the inherent structure-awareness of communities, we reduce the dependency on costly end-to-end training while improving inference accuracy. (b) Parameter-Free Structure Reconstruction: A tree-based sampler enables lightweight graph restructuring without trainable parameters. This avoids the computational overhead of gradient-based adjacency matrix optimization (e.g., nuclear norm regularization in ProGNN [10]) or probabilistic sampling (e.g., Gumbel-Softmax in NeuralSparse [37]).

### v5

@@ F.1 Topological Information Capturing Capability of the Encoding Tree
Lemma 1 (Locality and Additivity of Structural Entropy(Li and Pan, 2016)).
@@ F.2 Global information constraints of the low-level communities
the entropy term of the low-level community $\mathcal{C}^{\ell}_{\phi}$ becomes more localized (Li and Pan, 2016).
@@ F.3 The Error Bound Between Soft Labels and True Labels
The proof leverages three key lemmas derived from the Transformer architecture in (Vaswani et al., 2017).
@@
Let $t^{\star}_{\alpha}$ be the aggregated text of node $\alpha$ using threshold $\epsilon$ (Eq. 3). The raw soft label error $\Delta_{\alpha}^{\text{raw}}=|y^{cls}_{\alpha}-y_{\alpha}$— satisfies:
@@
the threshold $\epsilon$ filters neighbors with $w_{\alpha\beta}\geq\epsilon$ (Eq. 3 and Eq. 4). The multi-head attention reinforces local semantic consistency across subspaces, bounding the error proportionally to $(1-\epsilon)$ (Vaswani et al., 2017). ∎
@@
The residual structure $\text{LayerNorm}(x+\text{Sublayer}(x))$ suppresses feature drift, with $d_{\text{model}}$ as the hidden dimension (Vaswani et al., 2017):
@@
The learning rate design $lrate=d_{\text{model}}^{-0.5}\cdot\min(\cdot)$ indicates that increasing the model dimensionality necessitates a reduction in learning rate to preserve training stability. A larger $d_{\text{model}}$ enhances the model’s feature representation capacity, while the error decay rate per unit dimension scales proportionally with $d_{\text{model}}^{-0.5}$ (Vaswani et al., 2017). ∎
@@ Appendix G Datasets
We evaluate LLaTA and baselines on 11 widely adopted TAG datasets across multiple domains, including 4 citation networks (Chen et al., 2024), 1 knowledge network (Chen et al., 2024), 2 social networks (Huang et al., 2024), and 4 e-commerce networks (Yan et al., 2023). Among these, the Ratings and Child exhibit heterophily, while the remaining datasets follow homophily. The details of these TAG datasets are shown in Table 6. The description of the datasets for each domain is as follows:
@@
  • Citation Networks  (Chen et al., 2024). Cora, Citeseer, Pubmed and ArXiv are benchmark datasets of citation networks. Nodes represent paper, and edges represent citation relationships. The features are obtained through pre-trained language model, and labels denote their academic fields.
  • Knowledge Networks  (Chen et al., 2024; Mernyei and Cangea, 2020). WikiCS is a benchmark dataset of knowledge networks. Nodes represent articles in the field of computer science, and edges represent hyperlinks between these articles. The features are obtained through pre-trained language model, and labels denote different branches of computer science.
  • Social Networks  (Huang et al., 2024). Instagram and Reddit are benchmark datasets of social netowrks. Nodes represent users, and edges represent social relationhships. The features are obtained through pre-trained language model, and labels correspond to different types of users.
  • E-commerce Networks  (Yan et al., 2023). Ratings, Child, History and Photo are benchmark datasets of e-commerce networks. Nodes represent products, and edges represent relationships such as co-purchase or co-view. The features are obtained through pre-trained language model, and labels correspond to product categories or user ratings.
@@ Appendix H Baselines
In this section, we provide a brief description for each baseline used in the experiments. It includes 5 coupled GSL methods (Chen et al., 2020; Fatemi et al., 2021; Zhao et al., 2021; Wu et al., 2023; Zou et al., 2023), 6 decoupled GSL methods (Jin et al., 2020; Liu et al., 2022b; Li et al., 2022; Liu et al., 2022a; Nguyen et al., 2023; Bi et al., 2024), and 3 LLM-based GSL methods(Guo et al., 2024; Zhang et al., 2024). Based on this, we apply prevalent GNN (GCN (Kipf and Welling, 2016), GAT (Veličković et al., 2017), GraphSAGE (Hamilton et al., 2017)) and mainstream LLM-GNN (GLEM (Zhao et al., 2022), ENGINE (Zhu et al., 2024)) as the downstream backbone.
@@
  • GCN (Kipf and Welling, 2016). GCN is among the most widely adopted GNN architectures, as it introduces a first-order approximation of a localized spectral filter for graph-structured data.
@@
  • GAT (Veličković et al., 2017). GAT incorporates a self-attention mechanism to compute importance scores for different neighboring nodes, enabling more effective aggregation of neighborhood information.
@@
  • GraphSAGE (Hamilton et al., 2017). GraphSAGE [14] is an inductive framework that generates node embeddings by leveraging node features and using a sampling-based approach to aggregate information from the local neighborhood, enabling scalable representation learning.
@@
  • GLEM (Zhao et al., 2022). GLEM proposes a scalable framework that integrates LLMs with GNNs through a variational EM algorithm. By alternating between optimizing LMs for textual semantics and GNNs for structural information, GLEM enables mutual distillation without costly end-to-end training, achieving promising performance on large-scale text-attributed graphs.
@@
  • ENGINE (Zhu et al., 2024). ENGINE introduces an LLM-enhanced graph learning framework that combines LLMs with GNNs through a mutual integration mechanism. It leverages LLMs to enrich node representations with semantic knowledge while using GNNs to capture structural dependencies, and aligns the two via iterative refinement. This design improves both effectiveness and robustness on text-attributed graphs, showing promising performance across multiple benchmarks.
@@
  • IDGL (Chen et al., 2020). IDGL is a framework that learns a refined graph structure by leveraging cosine similarity of features and top-k threshold refinement, combining it with the original graph for joint optimization.
@@
  • SLAPS (Fatemi et al., 2021). SLAPS is a graph learning framework that uses MLP-based embeddings and kNN to build the graph. The adjacency matrix is symmetrized and normalized, and a self-supervised denoising autoencoder is used to update both the graph and model parameters.
@@
  • GAUGO (Zhao et al., 2021). GAUGO uses a graph auto-encoder to learn a new graph structure. It predicts edges based on node features, refines the edge probabilities with Gumbel sampling, and fuses the generated graph with the original one before training. Both the generated graph and model parameters are updated during optimization.
@@
  • HESGSL (Wu et al., 2023). HESGSL uses hierarchical embeddings to capture both local and global graph structures. It employs self-supervised learning with clustering and attention mechanisms to enhance feature representation and performance on downstream tasks.
@@
  • SEGSL (Zou et al., 2023). SEGSL constructs a graph by fusing a kNN graph with the original graph, leveraging structural entropy and encoding tree to refine edges and hierarchically extract community structures. The graph and model parameters are optimized jointly during training.
@@
  • ProGNN (Jin et al., 2020). ProGNN refines graph structures by enforcing low-rank, sparsity, and similarity constraints, jointly optimizing the learned graph and model parameters without predefined GSL bases or graph fusion.
@@
  • SUBLIME (Liu et al., 2022b). SUBLIME constructs GSL graphs using an anchor view (original graph) and a learner view. The learn view is initialized with kNN and optimized with parameterized or non-parameterized methods, followed by post-processing steps like top-k filtering and symmetrization. Contrastive learning between the two views refines the graph for downstream tasks.
@@
  • STABLE (Li et al., 2022). STABLE employs Graph Contrastive Learning (GCL) to generate robust graphs by perturbing node similarities and edges. Positive samples use slight perturbations, while negative samples use shuffled features. The graph refinement step applies a top-k filtering strategy on the node similarity matrix to retain helpful edges while removing adversarial ones.
@@
  • CoGSL (Liu et al., 2022a). CoGSL constructs GSL graphs by combining the original graph with additional views, such as adjacency or kNN graphs. Node embeddings are generated via GCNs, and connection probabilities are estimated and refined using InfoNCE loss to maximize mutual information. The final graph is obtained through early fusion and updated with model parameters.
@@
  • BORF (Nguyen et al., 2023). BORF is a curvature-based graph rewiring approach aimed at addressing over-smoothing and over-squashing in GNNs. It leverages Ollivier-Ricci curvature to identify edges with extreme curvature values—positive for over-smoothing and negative for over-squashing. This process preserves the graph’s topology while improving message-passing efficiency and enhancing GNN performance on downstream tasks.
@@
  • DHGR (Bi et al., 2024). DHGR enhances GNN performance on heterophily graphs by preprocessing the graph structure through rewiring. It systematically adds homophilic edges and removes heterophilic ones based on label and feature similarities, improving node classification accuracy, particularly for heterophily settings.
@@
  • GraphEdit (Guo et al., 2024). GraphEdit refines graph structures by integrating LLMs with a lightweight edge predictor to enhance graph representation learning. By leveraging instruction-tuning and training edge predictor, GraphEdit effectively denoises noisy connections and identifies implicit relationships among nodes.
@@
  • LLM4RGNN (Zhang et al., 2024). LLM4RGNN is a framework that enhances the adversarial robustness of GNNs using LLMs. It detects malicious edges and restores critical ones through a purification strategy, leveraging fine-tuned LLMs for edge prediction and correction. This approach effectively mitigates topology attacks, improving GNN performance across attack scenarios.
@@
  • LangGSL (Su et al., 2024). LangGSL presents a novel framework that integrates LLMs with GSL in an end-to-end manner. It simultaneously optimizes node features and graph connectivity by using LLMs to denoise textual attributes while GSL modules refine structural relations. This joint design enhances both semantic understanding and topological quality, leading to promising performance across diverse text-attributed graph datasets.
@@ Appendix I Experimental Setup
Current graph structure learning (GSL) methods face significant efficiency bottlenecks, particularly in large-scale or noisy graphs. Traditional approaches—including metric learning, probabilistic modeling, and direct optimization—often require computationally expensive pairwise similarity calculations (e.g., kernel functions or attention mechanisms) or iterative optimization of adjacency matrices(Zhu et al., 2021). These operations exhibit quadratic or higher complexity relative to node counts, limiting scalability. For LLM-based GSL Methods (e.g., GraphEdit (Guo et al., 2024), LLM4RGNN (Zhang et al., 2024)), efficiency issues are further exacerbated by the inherent overhead of large-model inference and fine-tuning. Empirical results reveal that these methods exceed 72 hours on larger datasets among the 11 benchmarks (As shown in Table 1), primarily due to their reliance on full-parameter tuning and dense graph operations.
Compared to existing paradigms, our framework achieves superior efficiency by: (a) Topology-Aware Context Enhancement: The Community-of-Thought mechanism provides high-quality topological context to guide LLM reasoning, eliminating the need for fine-tuning. By leveraging the inherent structure-awareness of communities, we reduce the dependency on costly end-to-end training while improving inference accuracy. (b) Parameter-Free Structure Reconstruction: A tree-based sampler enables lightweight graph restructuring without trainable parameters. This avoids the computational overhead of gradient-based adjacency matrix optimization (e.g., nuclear norm regularization in ProGNN (Jin et al., 2020)) or probabilistic sampling (e.g., Gumbel-Softmax in NeuralSparse (Zheng et al., 2020)).
@@
Appendix O Use of Large Language Models in paper writing
In preparing this manuscript, we made use of generative artificial intelligence (GenAI) tools, specifically GPT-4o and GPT-5, to assist with text refinement and polishing, as well as with the drafting and modification of auxiliary code snippets. These tools were employed solely to enhance clarity and readability and to streamline the presentation of supporting materials. Importantly, GenAI was not involved in the derivation of mathematical formulas, the design or implementation of core algorithms, or the development of key scientific insights. All theoretical analyses, algorithmic contributions, and experimental validations were carried out independently by the authors to ensure the originality and integrity of the work. Furthermore, all AI-assisted outputs were rigorously reviewed and verified by the authors to guarantee their accuracy and consistency with the scientific content, thereby upholding the reliability of the results presented.